# Enhancing Target-unspecific Tasks through a Features Matrix

### v4

@@ 2 Preliminaries
where $\operatorname{cos}(\cdot,\cdot)$ stands for cosine similarity, and $\tau$ represents a temperature parameter. In some learned text embedding methods, the method aligns with the CoOp methodology by setting the embeddings to be shared across different classes. During the inference phase, the prompts, along with the learned embeddings, can generate textual features that are used for classification purposes. By leveraging these learned embeddings, the model can produce more effective representations, leading to improved target-specific classification performance (Zhou et al., 2022b).
@@ 3 Methodology
3.1 Challenge and Motivation
As shown in Table 1, such as textual prompting (Zhou et al., 2022a, b; Yao et al., 2023a; Lu et al., 2022) and multi-modal prompting (Chen et al., ; Li et al., 2024a), have shown significant improvements in target-specific base classes. However, these methods often suffer from overfitting, leading to poor generalization in novel classes. To improve the novel performance of models, prompting regularization involves constraining the pre-trained and fine-tuning features in the text branch through computed loss. This constraint helps prevent forgetting general knowledge and ensures that the model retains essential information for downstream tasks. The objective of this design is to activate and maintain the pre-trained general knowledge, leveraging its remarkable generalization abilities to mitigate performance degradation when dealing with target-unspecific classes.
@@ 3.1 Challenge and Motivation
Method Design Base Novel
CLIP (ICML2021) Hand 69.34 74.22
CoCoOp (CVPR2022) Tune 80.47 71.69
ProDA (CVPR2022) Tune 81.56 72.30
PLOT (ICLR2023) Tune 81.24 72.98
ProGrad (ICCV2023) Tune 82.48 70.75
WCA (ICML2024) Tune 84.30 74.01
CoOp (IJCV2022) Tune 82.38 67.96
+ DePT (CVPR2024) Tune 83.66 71.82
+ Ours Tune + Hand 81.15 74.66
It is important to recognize that the fixed CLIP’s fundamental pre-trained characteristics demonstrate robust generalization capabilities. Recently, KgCoOp (Yao et al., 2023a) incorporates the constraint with $L1$ loss on fine-tuning feature with the pre-trained feature for the textual branch, avoiding the loss of CLIP’s original generalization capability. However, the novel performance of KgCoOp ($72.70\%$) is still lower than pre-trained CLIP ($74.22\%$(Radford et al., 2021) of Table 1. It may be attributed to the fact that this single hand-crafted prompting regularization with CLIP cannot fully release and excavate the diverse semantics for utilizing essential general knowledge (Khattak et al., 2023b).
@@ 3.2 Proposed Features Matrix (FM)
Table 2: Base-to-novel generalization. Compared with FAP (NIPS2024) and an easy-to-use DePT (CVPR2024) in detail, our method achieves consistent average performance improvement over different representative baselines (CoOp, CoCoOp, MaPLe, PromptSRC).
@@ 4.1 Benchmark Setting
FAP (Zhou et al., 2024) 72.00 93.40 90.15 65.61 70.13 85.89 23.70 67.88 45.98 45.12 68.99 65.68
Compared Methods. In three generalization experiments, we compare our method with CoOp (IJCV2022) (Zhou et al., 2022b), CoCoOp (CVPR2022) (Zhou et al., 2022a), MaPLe (CVPR2023) (Khattak et al., 2023a), PromptSRC (ICCV2023) (Khattak et al., 2023b), FAP (NIPS2024) (Zhou et al., 2024), and DePT (CVPR2024) (Zhang et al., 2024). The DePT is an easy-to-use method. We process frozen CLIP using a linear probe method. Comparison methods are based on the ViT-B/16 architecture for fair comparison.
@@ 4.2 Base-to-Novel Generalization Task
Discussion. In Table 2, our method demonstrates significant improvements on all $11$ datasets on HM. We found that our performance is higher than DePT (Zhang et al., 2024) (CVPR2024) when integrated with different methods. Compared to FAP (Zhou et al., 2024) (NIPS2024), our method significantly enhances the performance of base classes while also improving the accuracy on novel classes. In terms of average performance, our method based on PromptSRC (S) attains $85.70\%$ accuracy on the base classes, and $77.35\%$ accuracy on the novel classes. Our method, as an easy-to-use module, achieves performance improvements for various baseline frameworks, such as CoOp, CoCoOp, MaPLe (M), and PromptSRC (S).
@@ 4.3 Domain Generalization Task
We train our model using the ImageNet in 16 shots and test its performance on 4 different variants of the ImageNet. The training ImageNet is a non-generalizable setting in Table 3 (column-1). This experiment aims to validate the potential of our method in domain shifts (Fang et al., 2023).
@@ 6 Related Work
Prompt Tuning in VLMs. Prompt tuning (Zhang & Tian, 2025; Pan et al., 2024; Cao et al., 2025) is a commonly employed technique in the field of Natural Language Processing (NLP) for training on downstream tasks (Yu et al., 2025, 2024b, 2024a). Leveraging text prompts, which are instructions given to the language model component of VLMs, is a prevalent practice to improve task comprehension (Cui et al., 2025b). Full fine-tuning and linear probes are two commonly employed approaches for adapting VLMs to downstream tasks. The constraints of both methods have prompted research into innovative techniques influenced by prompt tuning within the realm of VLMs. CoOp (Zhou et al., 2022b) and CoCoOp (Zhou et al., 2022a) fine-tune the CLIP model for few-shot image recognition by optimizing a continuous set of embeddings within the textual branch. The image-conditional prompt utilized in CoCoOp significantly contributes to improving generalization to unseen classes. By conditioning prompts on visual features, CoCoOp (Zhou et al., 2022a) ensures that the language model focuses on pertinent visual information when making predictions. Moreover, some approaches (Yao et al., 2023b, a; Cui et al., 2024, 2025a) constrain the learnable prompts to contain the essential general knowledge. In addition, the approach (Zhang et al., 2024) introduces a generic and flexible approach to align its vision and language representations.
@@ 7 Future Works and Limitations
In the future, we plan to investigate the potential of our method in other tasks (Tian et al., 2024; Yang et al., 2023; Li et al., 2019; Yu et al., 2023; Li et al., 2024c; Zhang et al., 2022; Wang et al., 2021; Li et al., 2025a, c; Jia et al., 2025; Liu et al., 2023b, a; Li et al., 2024b) and downstream scenarios (Pan et al., 2025; Jin et al., 2025; Li et al., 2025b; Liu et al., 2024; Wu et al., 2024). Due to multiple processes of our method, the training speed is slower.

### v5

@@ 2 Preliminaries
where $\operatorname{cos}~{}(\cdot,\cdot)$ stands for cosine similarity, and $\tau$ denotes a temperature parameter. In some learned text embedding methods, the method aligns with the CoOp methodology by setting the embeddings to be shared across different classes. During the inference phase, the prompts, along with the learned embeddings, can generate textual features that are used for classification purposes. By leveraging these learned embeddings, the model can produce more effective representations, leading to improved target-specific classification performance (Zhou et al., 2022b).
@@ 3 Methodology
3.1 Current Challenge
As shown in Table 1, such as textual prompting (Zhou et al., 2022a, b; Yao et al., 2023a; Lu et al., 2022) and multi-modal prompting (Chen et al., ), have shown significant improvements in target-specific base classes. However, these methods often suffer from overfitting, leading to poor generalization in novel classes. To improve the novel performance of models, prompting regularization involves constraining the pre-trained and fine-tuning features in the text branch through computed loss. This constraint helps prevent forgetting general knowledge and ensures that the model retains essential information for downstream tasks. The objective of this design is to activate and maintain the pre-trained general knowledge, leveraging its remarkable generalization abilities to mitigate performance degradation when dealing with target-unspecific classes.
@@ 3.1 Current Challenge
Method Hand Features Base Novel
CLIP (ICML2021) Single 69.34 74.22
CoCoOp (CVPR2022) No 80.47 71.69
ProDA (CVPR2022) No 81.56 72.30
PLOT (ICLR2023) No 81.24 72.98
ProGrad (ICCV2023) No 82.48 70.75
CoOp (IJCV2022) No 82.38 67.96
+ DePT (CVPR2024) No 83.66 71.82
+ Kg (CVPR2023) Single 80.73 72.70
+ Ours Matrix 81.15 74.66
It is important to recognize that the fixed CLIP’s fundamental pre-trained characteristics demonstrate robust generalization capabilities. Recently, KgCoOp (Yao et al., 2023a) incorporates the constraint with $L1$ loss on fine-tuning feature with the pre-trained feature for the textual branch, avoiding the loss of CLIP’s original generalization capability.
Motivation. However, the novel class performance of hand-crafted-based KgCoOp ($72.70\%$) is still lower than pre-trained CLIP ($74.22\%$(Radford et al., 2021) of Table 1. It may be attributed to the fact that this single hand-crafted prompting regularization with CLIP cannot fully release and excavate the diverse semantics for utilizing essential general knowledge (Khattak et al., 2023b).
@@ 3.2 Proposed Features Matrix (FM)
Table 2: Base-to-novel generalization. In the base-to-novel generalization task, the datasets are divided into base and novel classes. The model is trained on the base classes in a 16-shot setting, and tested on both the base and novel classes across 11 different datasets. Compared with easy-to-use DePT (CVPR2024) in detail, our method achieves consistent average performance improvement over different representative baselines (CoOp, CoCoOp, MaPLe, PromptSRC).
@@ 4.1 Benchmark Setting
Compared Methods. In three generalization experiments, we compare our method with CoOp (IJCV2022) (Zhou et al., 2022b), CoCoOp (CVPR2022) (Zhou et al., 2022a), MaPLe (CVPR2023) (Khattak et al., 2023a), PromptSRC (ICCV2023) (Khattak et al., 2023b), and DePT (CVPR2024) (Zhang et al., 2024). The DePT is an easy-to-use method. We process frozen CLIP using a linear probe method. Comparison methods are based on the ViT-B/16 architecture for fair comparison.
@@ 4.2 Base-to-Novel Generalization Task
Discussion. In Table 2, our method demonstrates significant improvements on all $11$ datasets on HM. We found that our performance is higher than DePT (Zhang et al., 2024) (CVPR2024) when integrated with different methods. Compared to DePT, our method significantly enhances the performance of base classes while also improving the accuracy on novel classes. In terms of average performance, our method based on PromptSRC (S) attains $85.70\%$ accuracy on the base classes, and $77.35\%$ accuracy on the novel classes. Our method, as an easy-to-use module, achieves performance improvements for various baseline frameworks, such as CoOp, CoCoOp, MaPLe (M), and PromptSRC (S).
@@ 4.3 Domain Generalization Task
We train our model using the ImageNet in 16 shots and we leverage ImageNetA, ImageNet-R, ImageNet-Sketch, and ImageNetV2 to assess the model’s performance. The training ImageNet is a non-generalizable setting in Table 3 (column-1). This experiment aims to validate the potential of our method in domain shifts (Fang et al., 2023).
@@ 6 Related Work
Prompt Tuning in VLMs. Prompt tuning (Zhang & Tian, 2025; Pan et al., 2024; Cao et al., 2025; Zhou et al., 2024; Li et al., 2024a) is a commonly employed technique in the field of Natural Language Processing (NLP) for training on downstream tasks (Yu et al., 2025, 2024b, 2024a). Leveraging text prompts, which are instructions given to the language model component of VLMs, is a prevalent practice to improve task comprehension (Cui et al., 2025b). Full fine-tuning and linear probes are two commonly employed approaches for adapting VLMs to downstream tasks. The constraints of both methods have prompted research into innovative techniques influenced by prompt tuning within the realm of VLMs. CoOp (Zhou et al., 2022b) and CoCoOp (Zhou et al., 2022a) fine-tune the CLIP model for few-shot image recognition by optimizing a continuous set of embeddings within the textual branch. The image-conditional prompt utilized in CoCoOp significantly contributes to improving generalization to unseen classes. By conditioning prompts on visual features, CoCoOp (Zhou et al., 2022a) ensures that the language model focuses on pertinent visual information when making predictions. Moreover, some approaches (Yao et al., 2023b, a; Cui et al., 2024, 2025a) constrain the learnable prompts to contain the essential general knowledge. In addition, the approach (Zhang et al., 2024) introduces a generic and flexible approach to align its vision and language representations.
@@ 7 Future Works and Limitations
In the future, we plan to investigate the potential of our method in other tasks (Tian et al., 2024; Yang et al., 2023; Li et al., 2019; Yu et al., 2023; Li et al., 2024c; Zhang et al., 2022; Wang et al., 2021; Li et al., 2025a, c; Jia et al., 2025; Liu et al., 2023b, a; Li et al., 2024b) and downstream scenarios (Pan et al., 2025; Jin et al., 2025; Li et al., 2025b; Liu et al., 2024; Wu et al., 2024; Liu et al., 2023c, d). Due to multiple processes of our method, the training speed is slower. And, the ImageNet training in Table 4 is lower.